A trucks tires have the correct air pressure at room temperature (about 20 degrees Celsius) Explain why a tire is more likely to explode when the truck is driven in hot weather than cold weather
1
Expert's answer
2019-01-28T05:17:04-0500

According to Combined and Ideal Gas Laws

P1V1/T1 = P2V2/T2

If the volume of a tire is constant, then

P1/T1 =P2/T2

Consider the first situation when the weather is hot, for example the temperature of air is 40C.

P1/(273.15 + 20) = P2 /(273.15+ 40)

P2 = 1.068 * P1

The pressure inside the tire increases when the outside temperature increases. At a certain moment the pressure could become so high that the tire could explode .

Cosider the second situation when the weather is cold, for example the outside tempeture is T = -10C.

P1/(273.15+20) = P2/(273.15 -10)

P2 = 0.898*P1

The pressure inside the tire decreases as the outside temperature decreases. In the second situation ( cold weather) the tire will not explode .